# GreekBarBench: A Challenging Benchmark for Free-Text Legal Reasoning and Citations

### v2

@@ 5.2 Manual Evaluation by Legal Experts
GPT-4.1-mini 0.723 0.856 \$\$
GPT-4.1 0.807 0.855 \$\$\$
Gemini-2.0-F 0.751 0.794 \$
L-Krikri-8B 0.747 0.751
Gemma-3-27B 0.819 0.749
Gemini-2.0-L-F 0.695 0.708 \$
GPT-4.1-nano 0.542 0.372 \$
Table 4: Comparison of judges on GBB-JME, using Simple-Judge (Simple-J) and Span-Judge (Span-J). The cost per one million input tokens is denoted by: \$ (less than \$0.3), \$\$ (less than \$1), and \$\$\$ (less than \$3).
In this section we present the manual evaluations that we collected for GBB-JME (§ 5.3), our Judge Meta-Evaluation benchmark for assessing LLM-judges on GreekBarBench. We obtain ground truth evaluations (Facts, Cited Articles, Analysis scores on a scale of 1 to 10) from two expert legal annotators—licensed Greek lawyers with law degrees and practical experience. The annotators were compensated for their time and expertise. They evaluated five LLMs on 87 questions drawn from three exam sessions (2024-A, 2023-A, and 2023-B), resulting in a total of 1,305 annotated samples. The models evaluated on all three exams were Claude-3.7-Sonnet, OpenAI-o1, GPT-4o, and Gemini-2.0-Flash. For the 2024 exam, we included the open-source Llama-3.1-70B; however, due to its poor performance and generation failures on several questions, we replaced it with Deepseek-R1 for the 2023-A and 2023-B exams. Annotations were managed with the open-source platform doccano. [^8] https://doccano.prio.org
The results for the 2024 exam are shown in Figure 1. All LLMs except Llama-3.1-70B pass the exam. However, all LLMs lag behind the average human expert performance and the 95th percentile (‘top-5%’). The best-performing models are Claude-3.7-Sonnet (7.55) and OpenAI-o1 (7.52).
5.3 Judge Meta-Evaluation (GBB-JME)
We evaluate seven LLMs as judges on our benchmark for meta-evaluation, GBB-JME, using both the Simple-Judge prompt (§3.1) and the Span-Judge prompt with span-based rubrics (§3.2). The results are presented in Table 4. Generally, leading models (GPT-4.1, Gemini-2.0-Flash and GPT-4.1-mini) showed significant improvement when utilizing the Span-Judge prompt, while weaker models (Llama-Krikri-8B, Gemma-3-27B, Gemini-2.0-Lite-Flash, GPT-4.1-nano) struggled with its complexity. GPT-4.1-mini and GPT-4.1 achieved the best performance at 0.856 SPA, and 0.855 SPA, respectively. The competitive performance of the smaller model, aligns with observations reported elsewhere (Niklaus et al., 2025).
Figure 2: Comparison of closed and open-weight LLMs on GreekBarBench with GPT-4.1-mini Span-Judge.
Due to its strong performance and lower cost, we adopted GPT-4.1-mini as the judge for all subsequent evaluations. The total cost for all evaluations using GPT-4.1-mini was under \$60. The open-weight Gemma-3-27B model paired with the Simple-Judge prompt serves as a cost-effective alternative, but we also encourage researchers to evaluate new LLMs as judges on the publicly available GBB-JME benchmark.
Models Civil Law Criminal Law Commercial Law Public Law Lawyers’ Code Overall
f c a avg f c a avg f c a avg f c a avg f c a avg f c a avg
Top-5% - - - 9.00 - - - 10.00 - - - 10.00 - - - 9.20 - - - 9.18 - - - 8.87
Experts - - - 6.80 - - - 8.29 - - - 8.70 - - - 7.70 - - - 7.39 - - - 7.78
Gemini-2.5-F 8.8 8.4 8.4 8.53 8.4 7.9 8.5 8.28 8.6 8.0 8.3 8.27 8.7 8.0 8.1 8.28 8.9 8.5 8.5 8.62 8.7 8.2 8.4 8.40
GPT-4.1 8.8 8.1 8.4 8.44 8.5 8.2 8.2 8.28 8.4 8.1 8.3 8.27 8.6 7.7 8.1 8.14 8.8 8.2 8.4 8.48 8.6 8.0 8.3 8.32
Claude-3-7 8.5 7.2 7.4 7.72 8.2 6.9 7.0 7.37 7.6 6.9 7.4 7.31 8.6 7.2 7.6 7.79 8.5 8.2 8.2 8.29 8.3 7.3 7.5 7.71
GPT-4.1-mini 8.3 7.1 7.3 7.57 7.7 6.4 6.9 7.01 8.4 7.4 7.4 7.76 8.4 7.3 7.6 7.75 8.5 7.6 7.9 7.98 8.3 7.2 7.4 7.63
Gemma-3-27B 7.8 5.5 5.9 6.39 6.7 4.9 4.9 5.51 6.8 5.2 5.6 5.88 8.1 5.9 6.1 6.68 7.8 6.6 6.6 7.01 7.5 5.7 5.8 6.33
L-Krikri-8B 7.0 5.3 5.6 5.95 7.1 5.3 5.1 5.84 6.7 5.3 5.4 5.79 7.9 6.3 6.2 6.78 7.4 6.4 6.4 6.74 7.2 5.7 5.8 6.24
Table 5: Fine-grained comparison of proprietary and small open-weight LLMs on different legal areas: ‘Civil’, ‘Criminal’, ‘Commercial’, ‘Public’, ‘Lawyers’; and for different scoring dimensions: ‘Facts’ (f), ‘Cited Articles’ (c), ‘Analysis’ (a). ‘Experts’/‘Top-5%’ is the average/95th percentile score of candidate lawyers. Best LLM scores are in bold, failed scores are highlighted in red, and scores outperforming experts are highlighted in green.
@@ 5.4 Results on GreekBarBench
We conduct an extensive automatic evaluation of $13$ LLMs on GreekBarBench (Figure 2). We use GPT-4.1-mini as the LLM-Judge, employing the ‘Span LLM-Judge’ prompt (§ 3.2). The evaluation includes proprietary models such as GPT-4o, the GPT-4.1 family (GPT-4.1-mini, GPT-4.1-nano), Gemini-2.0-flash, and Claude-3.7-Sonnet (with reasoning disabled), along with the reasoning models OpenAI-o1 and Gemini-2.5-Flash. The open-weight models include the Gemma-3 family(Gemma-3-27B, Gemma-3-12B, and Gemma-3-4B), the specialized Greek model Llama-Krikri-8B-Instruct (Krikri-8B), and the reasoning model DeepSeek-R1.
@@
The experimental results (Figure 2) reveal that Gemini-2.5-Flash (8.4) and GPT-4.1 (8.32) demonstrate the strongest performance on GreekBarBench. They surpass the average legal expert score (7.78), though they still fall short of the 95th percentile (top-5%) of experts (8.87). OpenAI-o1 (7.77) and Claude-3.7-Sonnet (7.71), perform comparably to the average human expert (7.78). The fact that Gemini-2.5-Flash and OpenAI-o1, two reasoning models, are among the top performers, shows that leveraging inference-time reasoning is a key factor for this benchmark. The smallest models, GPT-4.1-nano (5.51) and Gemma-3-4B (4.41) are the only models that fail the exams (passing score: 6.00). The 8B Krikri model surpasses Gemma-12B and achieves performance comparable to the significantly larger Gemma-27B, highlighting the benefit of language-specific pretraining.
Table 5 presents a fine-grained comparison for four of the best proprietary LLMs (Gemini-2.5-Flash, GPT-4.1, Claude-3.7-Sonnet , GPT-4.1-mini) and two small open-weight LLMs (Gemma-3-27B, L-Krikri-8B), reporting their scores across five different legal areas (Civil, Criminal, Commercial, Public, Lawyers) and three scoring dimensions (Facts, Cited Articles, Analysis).
@@
The fine-grained comparison shows that LLMs exhibit consistent performance across all legal areas and scoring dimensions. Notably, in the areas of ‘Civil Law‘, ‘Public Law‘, and ‘Lawyers’ Code‘, four of the LLMs achieve scores surpassing the average human expert performance highlighted in green in Table 5. In contrast, the smaller open models, Gemma-3-27B and L-Krikri-8B (indicated by red), struggle in certain areas, failing to meet the passing grade threshold of 6.00. The second-best model, GPT-4.1, matches the top performer, Gemini-2.5-Flash, in ‘Criminal’ and ‘Commercial Law’, but Gemini-2.5-Flash achieves slightly higher scores in the remaining three areas. Interestingly, even though ‘Civil Law‘ seems to present the greatest challenge for human experts (average score of 6.80), this is not the case for LLMs.
Analyzing performance by dimension provides valuable insights into model capabilities. A key finding is that all models struggle most with the ‘Cited Articles‘ dimension. The higher overall scores of Gemini-2.5-Flash and GPT-4.1 are largely attributable to their stronger performance in the Cited Articles and Analysis dimensions, when compared to Claude-3.7-Sonnet and GPT-4.1-mini. Furthermore, Gemma-3-27B’s stronger performance compared to L-Krikri-8B is primarily attributable to its higher scores in the Facts dimension, while they get identical scores in the other dimensions.
@@ 7 Conclusions
In this work, we introduced GreekBarBench, a benchmark evaluating LLMs on legal questions requiring citations to statutory articles and case facts. We use a comprehensive scoring system and an LLM-judge for automatic evaluation. To ensure judge alignment with human experts, we developed an accompanying meta-evaluation benchmark (GBB-JME) using Soft-Pairwise Accuracy as the meta-metric. The results show that our span-based rubrics specifically designed for this benchmark improve the LLM-judges. The extensive evaluation of 13 LLMs on GreekBarBench revealed that Gemini-2.5-Flash and GPT-4.1 achieved the best performance, surpassing the typical human expert, but also highlighted areas for future improvement.
@@
Finally, the reported legal expert performance figures (average and 95th percentile) in our comparisons (§ 1, § 2) should be interpreted as illustrative baselines rather than rigorous head-to-head comparisons under identical evaluation conditions. This is due to inherent limitations in the available human data: detailed participant statistics are not available across all exam papers, and critically, the human scores were determined by the official Greek Bar Examination grading committee, not by our developed LLM-judge framework. Nevertheless, we believe these figures provide valuable intuition regarding the current performance gap between state-of-the-art LLMs and candidate lawyers.
@@ Acknowledgments
Appendix A Annotator Instructions

### v3

@@ 5.3 Judge Meta-Evaluation (GBB-JME)
GPT-4.1-mini 0.851 ± 0.008 0.862 ± 0.001 \$\$
GPT-4.1 0.835 ± 0.010 0.837 ± 0.005 \$\$\$
Gemini-2.0-F 0.745 ± 0.009 0.778 ± 0.011 \$
Gemma-3-27B 0.785 ± 0.015 0.733 ± 0.017
Krikri-8B 0.708 ± 0.004 0.638 ± 0.019
Gemini-2.0-LF 0.667 ± 0.016 0.695 ± 0.010 \$
GPT-4.1-nano 0.675 ± 0.006 0.416 ± 0.008 \$
Table 4: Comparison of LLMs-as-judges on GBB-JME, using Simple-Judge and Span-Judge. The cost per one million input tokens is denoted by: \$ (less than \$0.3), \$\$ (less than \$1), and \$\$\$ (less than \$3).
@@ 5.4 Results on GreekBarBench
We conduct an extensive automatic evaluation of $13$ LLMs on GreekBarBench (Figure 5). We use GPT-4.1-mini as the judge with the Span-Judge prompt (§ 3.2), reporting the Facts score, Articles score, Analysis score and Mean score with standard error from three runs. The evaluation includes proprietary models such as GPT-4o, the GPT-4.1 family (GPT-4.1, GPT-4.1-mini, GPT-4.1-nano), Gemini-2.0-Flash, and Claude-3.7-Sonnet (reasoning disabled), along with the reasoning models OpenAI-o1 and Gemini-2.5-Flash. The open-weight models include the Gemma-3 family (27B, 12B, and 4B parameters), the specialized Greek model Llama-Krikri-8B-Instruct (Krikri-8B), and the reasoning model DeepSeek-R1.
Model Facts Articles Analysis Mean
Gemini-2.5-F 8.62 8.16 8.36 8.38 ± 0.02
GPT-4.1 8.65 8.05 8.34 8.35 ± 0.02
OpenAI-o1 8.24 7.55 7.55 7.78 ± 0.01
Claude-3.7-S. 8.34 7.37 7.60 7.77 ± 0.04
GPT-4.1-mini 8.28 7.18 7.40 7.62 ± 0.01
Gemini-2.0-F 8.28 7.13 7.12 7.51 ± 0.03
GPT-4o 8.15 7.10 7.18 7.48 ± 0.03
Deepseek-R1 7.65 6.48 6.55 6.90 ± 0.02
Gemma-3-27B 7.46 5.63 5.79 6.29 ± 0.02
Krikri-8B 7.20 5.65 5.74 6.20 ± 0.03
Gemma-3-12B 7.25 5.37 5.62 6.08 ± 0.04
PASS score 6.00 6.00 6.00 6.00 ± 0.00
GPT-4.1-nano 6.88 4.69 4.79 5.45 ± 0.03
Gemma-3-4B 5.53 3.76 3.95 4.42 ± 0.00
Table 5: Comparison of closed and open-weight LLMs on GreekBarBench with GPT-4.1-mini Span-Judge on three runs. The scores reported are ‘Facts’, ‘Articles’, ‘Analysis’, and ‘Mean’ score with standard error. The best scores are shown in bold, while failed scores are highlighted in red.
@@
Gemini-2.5-Flash (8.38) and GPT-4.1 (8.35) achieved the highest scores on GreekBarBench (Table 5). These scores are not directly comparable to the human expert scores (average: 7.78, top-5%: 8.87) in Figure 1, as the models were evaluated by an LLM-judge on a larger dataset (310 vs. 22 questions). The smallest models, GPT-4.1-nano (5.45) and Gemma-3-4B (4.42) are the only models that fail the exams (passing score: 6.00). The open-weight Krikri-8B model (6.20) surpasses Gemma-3-12B (6.08) and achieves performance comparable to the significantly larger Gemma-3-27B (6.29), highlighting the benefit of language-specific pretraining.
Analyzing the scoring dimensions provides valuable insights into model capabilities. A key finding is that all models except for Gemini-2.5-Flash and GPT-4.1 struggle most with the ‘Articles‘ and the ‘Analysis‘ dimensions and that’s what separates them from the others.
@@
Table 6 presents a fine-grained comparison for six LLMs, reporting their mean score across five different legal areas (Civil, Criminal, Commercial, Public, Lawyers) and their average (Avg). The results show that LLMs exhibit consistent performance across all legal areas. Notably, the smaller open models, Gemma-3-27B and Krikri-8B, struggle in certain areas, failing to meet the passing grade threshold of 6.00 (indicated by red). The second-best model, GPT-4.1, matches the top performer, Gemini-2.5-Flash, in ‘Criminal’ and ‘Commercial Law’, but Gemini-2.5-Flash achieves slightly higher scores in the remaining three areas.
Model Civil Crim Comm Public Lawyer Avg
Gem-2.5 8.53 8.28 8.27 8.28 8.62 8.38
GPT-4.1 8.44 8.28 8.27 8.14 8.48 8.35
Claude-3.7-S 7.72 7.37 7.31 7.79 8.29 7.77
GPT-4.1-mini 7.57 7.01 7.76 7.75 7.98 7.62
Gem-27B 6.39 5.51 5.88 6.68 7.01 6.29
Krikri-8B 5.95 5.84 5.79 6.78 6.74 6.20
Table 6: Comparison of model performance on different legal areas: ‘civil’, ‘criminal’, ‘commercial’, ‘public’, ‘lawyer’ and the average score (‘Avg’).
5.5 Chapter-based Legal Context
Model no context chapter context oracle context
f c a avg f c a avg f c a avg
Gemini-2.5-Flash 8.2 6.1 6.8 7.02 8.6 8.2 8.4 8.38 8.9 8.8 8.5 8.73
GPT-4.1 7.8 5.8 6.4 6.63 8.7 8.1 8.3 8.35 8.7 8.9 8.6 8.73
GPT-4.1-mini 7.5 4.4 5.4 5.73 8.3 7.2 7.4 7.62 8.7 8.5 8.1 8.44
Gemini-2.0-Flash 7.7 5.0 5.5 6.05 8.3 7.1 7.1 7.51 8.6 8.2 7.8 8.19
Gemma-3-27B 7.1 4.1 4.9 5.37 7.5 5.6 5.8 6.29 8.5 8.0 7.7 8.08
Table 7: Comparison with different context settings. Detailed scores for facts (f), cited articles (c), analysis (a) and mean/average (avg) are provided. The no-context setting makes the task much more challenging (most models barely get a passing score, which is above 6.00). The oracle context setting allows for higher scores, because the models make fewer mistakes in the cited articles score.
To investigate the impact of legal context, we experimented with three settings, with results presented in Table 7. The “no-context” setting provided no legal corpus, forcing models to rely solely on their parametric knowledge. The “chapter” setting, our primary setup, provided relevant articles alongside distractors, as described in Section 2.4. The “oracle” setting provided only the exact correct articles. In the no-context setting, all models except Gemini-2.5-Flash performed poorly, underscoring their limited internal knowledge of Greek legislation. Conversely, the oracle setting enabled every model to achieve excellent scores, establishing that access to the correct legal statutes is the most critical factor for success in this task.
As expected, providing more accurate legal context—progressing from the no-context to the chapters and oracle settings—dramatically improved the cited articles score for all models. The facts and analysis scores also increased slightly, showing that access to the right articles improves the model’s reasoning.
Notably, Gemini-2.0’s initial lead over GPT-4.1-mini in the no-context setting (6.05 vs. 5.73) was overturned in the chapter (7.62 vs. 7.51) and oracle (8.19 vs 8.44) settings. This divergence demonstrates that strong parametric knowledge and the ability to process long context are separate skills, as shown in (Liu et al., 2024). This highlights the necessity of evaluating both abilities independently.
6 Manual Error Analysis
To supplement our quantitative metrics, we conducted a manual error analysis to identify common failure modes. We analyzed the 10 worst-scoring responses from five key models. Human experts, guided by the scores and explanations of the LLM-judge, categorized the errors to reveal the most significant weaknesses. We focused on critical errors that led to an incorrect answer. The types of errors are the following:
Facts:
Errors include hallucinations (‘Hall.’), inventing facts not present in the provided context, and inaccuracies (‘Inac.’), misinterpreting or omitting key facts from the context.
Cited articles:
Errors include inapplicable articles (‘Inap.’), citing legal articles not relevant to the case, and missing articles (‘Missing’), failing to identify essential articles for the analysis.
Analysis:
Errors include an invalid argument (‘Argum.’), a flawed argument from misunderstood principles or fallacious reasoning, and a generic answer (‘Generic’), providing a general statement instead of an analysis tailored to the case facts.
Our analysis, while based on a small sample, highlights three primary areas of weakness: citing legal articles, understanding case facts, and constructing valid arguments. The most frequent errors related to failing to cite the correct legal articles (23). This supports our finding (§5.5) that correct article retrieval is essential. Factual (20) and analysis (17) errors were also prevalent across all models. Although rare, critical fact hallucinations (3) appeared even in top-performing models, highlighting the risks of deploying these systems without human oversight.
Model Facts Articles Analysis
Hall.
Inac.
Inap.
Missed
Argum.
Generic
Gemini-2.5-F 1 3 1 0 3 1
GPT-4.1 1 2 3 5 2 1
Claude-3.7-S 0 3 0 3 5 0
GPT-4.1-mini 0 4 1 3 4 0
Gemma-3-27B 1 5 0 7 1 0
Total 20 23 17
Table 8: Manual analysis of error types across different models. Errors are grouped into three main categories: Facts, Articles, and Analysis; and two subcategories for each one: hallucinations (‘Hall.’) and inaccuracies (‘Inac.’) for facts, inapplicable (‘Inap.’) and missed (‘Missed’) articles, and invalid argument (‘Argum.’) and generic answer (‘Generic’) for the analysis.
A key differentiator between human and model performance was the inability of LLMs to grasp subtle, legally significant context. For instance, all LLMs failed to infer that the phrase “they had differences” could indicate a motive for premeditation, which was obvious to human experts. This failure to interpret nuanced language was a common weakness, showing that significant improvement is needed for reliable use in real-world applications.
@@ 8 Conclusions
In this work, we introduced GreekBarBench, a benchmark evaluating LLMs on legal questions requiring citations to statutory articles and case facts. To enable robust and scalable assessment, we developed a comprehensive evaluation framework based on LLM-as-a-judge, which we validated against human experts using our GBB-JME meta-evaluation benchmark. Our results show that custom span-based rubrics significantly improve judge alignment. The extensive evaluation of 13 LLMs on GreekBarBench revealed that Gemini-2.5-Flash and GPT-4.1 achieved the best performance. Our primary finding is that success on this task is critically dependent on the quality of the provided legal context and the ability to identify the correct statutory articles within it.
@@ Acknowledgments
Appendix A LLMs-as-judges variants
Model GPT-4.1-mini Gemma-3-27B Delta
Span Simple
Gemini-2.5-Flash 8.38 8.53 +0.15
GPT-4.1 8.35 8.38 +0.03
OpenAI-o1 7.78 8.05 +0.27
Claude-3-7-Sonnet 7.77 8.08 +0.32
GPT-4.1-mini 7.62 7.87 +0.25
Gemini-2.0-flash 7.51 7.88 +0.37
GPT-4o 7.48 7.75 +0.27
Deepseek-R1 6.90 7.21 +0.32
Gemma-3-27b 6.29 6.84 +0.55
Krikri-8B 6.20 6.80 +0.60
Gemma-3-12b 6.08 6.56 +0.48
GPT-4.1-nano 5.45 5.77 +0.31
Gemma-3-4b 4.42 5.03 +0.61
Table 9: Comparison of scores from two LLM-as-judge configurations: our primary judge (GPT-4.1-mini with Span-Judge) versus an alternative (Gemma-3-27B with Simple-Judge). The table is sorted by the GPT-4.1-mini scores. Scores from Gemma-3-27B are highlighted in bold where they alter the ranking of a model relative to its neighbors.
Table 9 presents a comparative analysis to assess the robustness of our evaluation framework. We compare the scores from our primary judge, GPT-4.1-mini with the Span-Judge prompt, against those from an alternative setup using Gemma-3-27B with a simpler Simple-Judge prompt.
Our analysis reveals two key findings. First, the overall model rankings are highly consistent across both judges, as evidenced by a strong SPA agreement of 0.92. This indicates that our top-level conclusions about model performance are robust to the choice of the judge.
Second, despite the high rank correlation, we observe a significant and systematic scoring bias. Gemma-3-27B acts as a more lenient judge, assigning scores that are, on average, 0.35 points higher than GPT-4.1-mini. Crucially, this bias is not uniform across all models. As shown in the Delta column, the score inflation is most pronounced for lower-capability models (e.g., +0.61 for Gemma-3-4b and +0.60 for Krikri-8B) and is minimal for top-tier models. The exceptionally small delta for GPT-4.1 (+0.03) is particularly noteworthy. This suggests that the judges disagree on its proximity to the top-ranked Gemini-2.5-Flash. The low delta could indicate a potential " “family bias,” where the GPT-4.1-mini judge evaluates its sibling model more favorably, rating it closer to the top performer than the external Gemma judge does.
This non-uniform bias leads to minor but notable disagreements in the fine-grained ordering of similarly-performing models. For example, the Gemma judge awards a disproportionately higher score to Claude-3-7-Sonnet and Gemini-2.0-Flash, causing them to overtake their immediate competitors in the ranking, as highlighted in bold. This experiment underscores that while overall rankings are stable, the absolute scores and the precise ordering of closely-ranked models can be sensitive to the specific judge and prompting method used.
Appendix B Contamination
Model Knowledge cutoff GreekBarBench Semi-private Delta
Gemini-2.5-Flash 2025-01-01 8.40 8.04 – 0.36
GPT-4.1 2024-06-01 8.32 7.65 – 0.67
OpenAI-o1 2023-09-01 7.77 7.51 – 0.26
Claude-3.7-Sonnet 2024-09-30 7.71 7.51 – 0.20
GPT-4.1-mini 2024-06-01 7.63 7.09 – 0.54
Gemma-3-27B 2024-08-30 6.33 5.51 – 0.82
Krikri-8B 2025-02-09 6.24 5.56 – 0.68
Semi-private set cutoff: 2024-10-03
Table 10: We report the overall performance of LLMs on the original GreekBarBench and on the semi-private set. Delta is the performance difference between the two sets. The models highlighted in bold could potentially have contaminated data and still trick the test, because their cutoff date is after the semi-private set’s date of publication.
Table 10 compares the performance of several LLMs on the public GreekBarBench and its ‘semi-private’ version. A significant drop in performance on the ‘semi-private’ set, indicated by a large negative delta, would suggest contamination. The results show no clear evidence of such contamination. The observed performance drops are minor for all models (under a 1.0-point delta), suggesting the ‘semi-private’ set is simply more difficult. The largest deltas are also associated with the lowest-performing models (Gemma-3-27B and Krikri-8B), which contradicts the expected score inflation.
The exceptions to this analysis are Gemini-2.5-Flash and Krikri-8B, whose late data cutoffs allow for potential data exposure to remain undetected. However, the concern for Krikri-8B is mitigated by its very low performance, which is inconsistent with contamination, as mentioned before. Consequently, only the results for Gemini-2.5-Flash are inconclusive. Our final assessment is that most models appear uncontaminated, while the status of Gemini-2.5-Flash cannot be determined.
Appendix C Annotator Instructions